OpenStudy (anonymous):

Check my last two answers please? (-10^9-4m^8-12m^6)/2m^4 a. -5m^5-2m^4-12m^6 b.-5m^5-2m^4-6m^2 c. -5m^9-2m^8-6m^6 - my answer d. -5m^5-4m^8-12m^2 Divide s^2+6s/s^2+10s+24 divide by s-4/s+4 a. s/s-4 b.s+6/s-4 c.s+6/s^2-4s d.s-4/s - my answer

OpenStudy (anonymous):

@Loser66 sorry to bother you again, could you just do me a favor and check these? I am unsure about these two

OpenStudy (loser66):

nope, redo the first one, you just factor 2m^4 out from the numerator and then cancel it with the denominator. Done

OpenStudy (anonymous):

so the first one would be b?

OpenStudy (loser66):

yup

OpenStudy (anonymous):

Thanks! and what about the 2nd one?

OpenStudy (loser66):

it is unclear to me, let re-write \[\Large \dfrac{\frac{s^2+6}{s^2+10s+24}}{\frac{s-4}{s+4}}\] right?

OpenStudy (anonymous):

no its this

OpenStudy (loser66):

oh, so your answer should be flipped, numerator becomes denominator and vice versa

OpenStudy (anonymous):

ok so a?

OpenStudy (loser66):

yup

OpenStudy (anonymous):

thank you!!

OpenStudy (loser66):

np